# Multi-Subarray FD-RIS Enhanced Multi-user Wireless Networks: With Joint Distance-Angle Beamforming

### v1

@@ III-A System Model
In this section, the efficacy of the multi-subarray FD-RIS is verified by considering the communication setup illustrated in Fig. 1. This system comprises a BS equipped with $N_{\mathrm{t}}$ antennas, $K$ users each equipped with a single antenna, and a multi-subarray FD-RIS configured similarly to the FD-RIS described in section II. It is assumed that the direct connection between the BS and users be blocked by buildings or trees, a situation frequently encountered in urban communication scenarios. Due to the existence of the LoS links between the BS and the FD-RIS, as well as between the FD-RIS and the users, the Rician channel model is utilized to represent the connections between the BS and the FD-RIS, i.e., $\mathbf{H}_{\mathrm{BR}}$, and also between the FD-RIS and the users, i.e., $\mathbf{h}_{\mathrm{r}k}$. In particular, channel $\mathbf{H}_{\mathrm{BR}}$ can be expressed as
$\mathbf{H}_{\mathrm{BR}}=\zeta(d_{\mathrm{br}})\left(\sqrt{\frac{\beta}{1+\beta}}\mathbf{H}_{\mathrm{BR}}^{\mathrm{LoS}}+\sqrt{\frac{1}{1+\beta}}\mathbf{H}_{\mathrm{BR}}^{\mathrm{NLoS}}\right),$ (10)
where $\beta$ is the Rician factor. $\mathbf{H}_{\mathrm{BR}}^{\mathrm{NLoS}}$ denotes the non line-of-sight (NLoS) component, and each element of $\mathbf{H}_{\mathrm{BR}}^{\mathrm{NLoS}}$ is independent identically distributed (i.i.d.) and follows complex Gaussian distribution with zero mean and unit variance. According to the signal processing model of the FD-RIS in section II, the LoS component, i.e., $\mathbf{H}_{\mathrm{BR}}^{\mathrm{LoS}}$, can be given as
@@
  • $\mathbf{a}_{\mathrm{R}}=\left[1,\cdots,e^{-j2\pi\frac{\Gamma^{\mathrm{br}}_{i_{z},i_{y}}}{\lambda}},\cdots,e^{-j2\pi\frac{\Gamma^{\mathrm{br}}_{I_{z},I_{y}}}{\lambda}}\right]^{T},$
  • $\mathbf{a}_{\mathrm{B}}=\Big{[}1,\cdots,e^{-j2\pi\frac{(n_{\mathrm{t}}-1)d\sin(\theta_{\mathrm{br}})\sin(\varphi_{\mathrm{br}})}{\lambda}},\cdots,\\ e^{-j2\pi\frac{(N_{\mathrm{t}}-1)d\sin(\theta_{\mathrm{br}})\sin(\varphi_{\mathrm{br}})}{\lambda}}\Big{]}^{T}$.
@@
$\mathbf{h}_{\mathrm{r}k}=\zeta(d_{\mathrm{r}k})\left(\sqrt{\frac{\beta}{1+\beta}}\mathbf{h}_{\mathrm{r}k}^{\mathrm{LoS}}+\sqrt{\frac{1}{1+\beta}}\mathbf{h}_{\mathrm{r}k}^{\mathrm{NLoS}}\right),$ (12)
where $\zeta(d_{\mathrm{r}k})$ denotes the large-scale path loss with $d_{\mathrm{r}k}$ being the distance between the FD-RIS and the $k$-th user. $\mathbf{h}_{\mathrm{r}k}^{\mathrm{NLoS}}$ is the NLoS component, and each element of $\mathbf{h}_{\mathrm{r}k}^{\mathrm{NLoS}}$ is i.i.d. and follows complex Gaussian distribution with zero mean and unit variance. $\mathbf{h}_{\mathrm{r}k}^{\mathrm{LoS}}$ represents the LoS component, which is given by
@@
  • $y_{k}=\mathbf{h}_{\mathrm{r}k}^{T}\tilde{\boldsymbol{\Theta}}\boldsymbol{\Theta}_{k}\mathbf{H}_{\mathrm{BR}}\mathbf{w}_{k}s_{k}+\sum_{j\neq k}^{K}\mathbf{h}_{\mathrm{r}k}^{T}\tilde{\boldsymbol{\Theta}}\boldsymbol{\Theta}_{k}\mathbf{H}_{\mathrm{BR}}\mathbf{w}_{j}s_{j}+n_{k},$
@@ III-B Problem Formulation
where $\omega_{k},~{}k\in\mathcal{K},$ is the weighted factor of the $k$-th user, which can be utilized to control the priority of the $k$-th user and the fairness among the users. $f_{\min}$ and $f_{\max}$ denote the lower and upper bounds of the modulation frequency ${f}_{l},$ for $l\in\mathcal{L}$. $P_{\mathrm{tmax}}$ is the maximum power budget of the BS. Actually, it is challenging to address this optimization problem, since the non-convexity of the objective function with respect to (w.r.t.) $\{\mathbf{w}_{k}\}_{k=1}^{K}$, $\mathbf{f}$ and $\boldsymbol{\tau}$. To handle this problem, the MMSE method [26] is first utilized to transform the non-convex objective function. Then, the alternating strategy is leveraged to divide the transformed optimization problem into three subproblems, which are respectively for the design of active beamforming, time-delay variables, and modulation frequency variables. Specifically, leveraging the MMSE method and introducing auxiliary variables $W_{k}$ and $\mu_{k}$, the communication rate $R_{k},~{}k\in\mathcal{K},$ can be converted as
@@
$\begin{cases}W_{k}^{\mathrm{opt}}=E_{k}^{-1},&\\ u_{k}^{\mathrm{opt}}=\frac{\mathbf{h}_{\mathrm{r}k}^{T}\tilde{\boldsymbol{\Theta}}\boldsymbol{\Theta}_{k}\mathbf{H}_{\mathrm{BR}}\mathbf{w}_{k}}{\sum_{j=1}^{K}\left|\mathbf{h}_{\mathrm{r}k}^{T}\tilde{\boldsymbol{\Theta}}\boldsymbol{\Theta}_{k}\mathbf{H}_{\mathrm{BR}}\mathbf{w}_{j}\right|^{2}+\sigma_{k}^{2}}.&\end{cases}$ (17)
The analysis above indicates that the lower bound of $R_{k}$ in the $(q+1)$-th iteration can be obtained with the given $W_{k}^{(q)}$ and $u_{k}^{(q)}$. Note that the values of $W_{k}^{(q)}$ and $u_{k}^{(q)}$ are calculated by utilizing the optimal expressions in (17) with the achieved active beamforming, time delays and modulation frequency solutions in the $q$-th iteration. Hence, the lower bound of $R_{k}$ in the $(q+1)$-th iteration can be expressed as $R_{k}\geq\tilde{R}_{k}=\frac{\ln(W_{k}^{(q)})-W_{k}^{(q)}E_{k}+1}{\ln 2}.$ Thus, in $(q+1)$-th iteration, the original optimization problem can be transformed as
@@ IV-A Active beamforming design
[…]
@@ IV-B Passive Beamforming Design for Time Delay
From the expression of $\tilde{\theta}_{i}=e^{-j2\pi g{f}_{l}\tau_{i}},~{}i\in\mathscr{I}$, it is worth noting that regardless the value of ${f}_{l},l\in\mathcal{L}$, each element in the FD-RIS has the capability to induce a specific phase-shift on the reflected harmonic signal by introducing an appropriate time delay. Therefore, we can initially design the reflection coefficients generated by the time delays in a holistic manner. In other words, we are directed towards the holistic design of $\tilde{\theta}_{i}=e^{-j2\pi g{f}_{l}\tau_{i}},~{}i\in\mathscr{I}$ rather than concentrating on the design of the $\tau_{i},~{}i\in\mathscr{I}$. Subsequently, it is straightforward to obtain the corresponding time delays $\tau_{i},~{}i\in\mathscr{I}$ with given ${f}_{l},~{}l\in\mathcal{L}$ based on the solved reflection coefficients $\tilde{\theta}_{i},~{}i\in\mathscr{I}$.
@@
[…]
@@
  • $\boldsymbol{\vartheta}^{*}=\operatorname{diag}(\tilde{\boldsymbol{\Theta}})$, $\mathbf{B}=\sum\limits_{k=1}^{K}\frac{\omega_{k}{W}_{k}^{(q)}\mathbf{B}_{k}}{\ln 2}$, $\mathbf{b}=\sum\limits_{k=1}^{K}\frac{\omega_{k}{W}_{k}^{(q)}\mathbf{b}_{k}}{\ln 2}$,
@@
[…]
@@
where $\mathscr{P}(\cdot)$ is the projection operator, which maps the updated $\boldsymbol{\vartheta}$ to the feasible region $\mathcal{S}$ by element-wise retraction, i.e., $\frac{(\boldsymbol{\vartheta})_{i}}{|(\boldsymbol{\vartheta})_{i}|},~{}i\in\mathscr{I}$. $\eta^{(e)}$ denotes the Armijo step in the $(e+1)$-th iteration. The details for solving the optimization problem (IV-B) utilizing the Riemannian conjugate gradient algorithm are presented in Algorithm 2.
@@
7: Output: the optimal $\boldsymbol{\vartheta}^{\mathrm{opt}}$ and the corresponding optimal
 time delay $\boldsymbol{\tau}^{\mathrm{opt}}$.
@@ IV-C Passive Beamforming Design for Modulation Frequency
  • $u_{l}^{(z)}=\begin{cases}{f}_{l}^{(z-1)}+f_{\max}-f_{\min},&z\leq 2,\\ {f}_{l}^{(z-1)}+\gamma^{(z)}\left(u_{l}^{(z-1)}-{f}_{l}^{(z-2)}\right),&z\geq 3,\\ \end{cases}$ denotes the upper asymptote in the $z$-th iteration.
  • $o_{l}^{(z)}=\begin{cases}{f}_{l}^{(z-1)}-f_{\max}+f_{\min},&z\leq 2,\\ {f}_{l}^{(z-1)}-\gamma^{(z)}\left({f}_{l}^{(z-2)}-o_{l}^{(z-1)}\right),&z\geq 3,\\ \end{cases}$ denotes the lower asymptote in the $z$-th iteration.
@@
  • $\rho^{(z)}=\frac{1}{10L}\operatorname{abs}\left(\frac{\partial{\color[rgb]{0.00,0.00,1.00}\tilde{g}}}{\partial\mathbf{f}}\left(\mathbf{f}^{(z-1)}\right)\right)^{T}\left(\mathbf{f}_{\max}-\mathbf{f}_{\min}\right)$ is the conservative factor, which can control the conservative property of the MMA convex approximation.
@@
In addition, the first-order gradients of the objective function w.r.t. the optimized variable, i.e., $\frac{\partial\tilde{g}}{\partial\mathbf{f}}$, can be derived as
@@
with
  • $\mathcal{I}_{v}\left(\frac{\partial\boldsymbol{\vartheta}_{k}}{\partial{f}_{l}}\right)=\begin{cases}A_{0}j2\pi g\left(\frac{d^{\mathrm{r}k}_{v}-t}{c}\right)\times\\ e^{j2\pi g{f}_{l}\left(\frac{d^{\mathrm{r}k}_{v}-t}{c}\right)+\phi_{0}},&v\in\mathcal{V},\\ 0,&otherwise,\end{cases}$
where we have $\mathcal{V}\triangleq\{(r-1)SMN+{\color[rgb]{0.00,0.00,1.00}\tilde{s}}N+(m-1)SN+n\}_{m\in\mathcal{M},n\in\mathcal{N}}$ with $r=\lceil l/S\rceil,{\color[rgb]{0.00,0.00,1.00}\tilde{s}}=\operatorname{mod}(l-1,S)$. Note that $d^{\mathrm{r}k}_{v}$ denotes the distance between the $v$-th element of the FD-RIS and user $k$, which can be mapped as $d^{\mathrm{r}k}_{i_{z},i_{y}}$ easily.
@@
$\frac{\partial\widehat{g}}{\partial{f}_{l}}(\mathbf{f}|\mathbf{f}^{(z-1)})=\frac{q_{l}^{(z)}}{\left(u_{l}^{(z)}-{f}_{l}\right)^{2}}-\frac{q_{l}^{(z)}}{\left({f}_{l}-o_{l}^{(z)}\right)^{2}}.$ (38)
@@
In order to address this issue, an additional inner loop iteration is implemented to fine-tune the conservative factor $\rho$ for a more conservative convex approximation. This adjustment aims to satisfy the condition $\widehat{g}(\mathbf{f}^{(z)}|\mathbf{f}^{(z-1)})\geq\tilde{g}(\mathbf{f}^{(z)})$. Specifically, if the solution of the MMA subproblem in the $z$-th outer loop iteration does not meet the third condition, we will employ the inner loop iteration to progressively enhance the conservative factor until the third condition is met. In the $(x+1)$-th inner loop iteration, the update rule of the conservative factor can be given by

### v2

@@ III-A System Model
In this section, the efficacy of the multi-subarray FD-RIS is verified by considering the communication setup illustrated in Fig. 2. This system comprises a BS equipped with a uniform linear array (ULA) of $N_{\mathrm{t}}$ antennas, $K$ users each equipped with a single antenna, and a multi-subarray FD-RIS  [^3] The modulation frequencies applied to different subarrays do not alter the inherent response of the RIS. Because the RIS’s responses for the incident signal depend solely on its carrier frequency. As long as the carrier frequency remains unchanged, the RIS responses under different states with respect to the incident signal, will not vary. It should be emphasized that the modulation frequencies in the RIS influence only the speed of phase-shift variation applied to the incident signal. configured similarly to the FD-RIS described in section II. It is assumed that the direct connection between the BS and users be blocked by buildings or trees, a situation frequently encountered in urban communication scenarios. Due to the existence of the LoS links between the BS and the FD-RIS, as well as between the FD-RIS and the users, the Rician channel model is utilized to represent the connections between the BS and the FD-RIS, i.e., $\mathbf{H}_{\mathrm{BR}}$, and also between the FD-RIS and the users, i.e., $\mathbf{h}_{\mathrm{r}k}$. In particular, channel $\mathbf{H}_{\mathrm{BR}}$ can be expressed as
$\mathbf{H}_{\mathrm{BR}}=\zeta(d_{\mathrm{br}})\left(\sqrt{\frac{\beta_{1}}{1+\beta}}\mathbf{H}_{\mathrm{BR}}^{\mathrm{LoS}}+\sqrt{\frac{1}{1+\beta_{1}}}\mathbf{H}_{\mathrm{BR}}^{\mathrm{NLoS}}\right),$ (13)
where $\beta_{1}$ is the Rician factor. $\mathbf{H}_{\mathrm{BR}}^{\mathrm{NLoS}}$ denotes the non line-of-sight (NLoS) component, and each element of $\mathbf{H}_{\mathrm{BR}}^{\mathrm{NLoS}}$ is independent identically distributed (i.i.d.) and follows complex Gaussian distribution with zero mean and unit variance. According to the signal processing model of the FD-RIS in section II, the LoS component, i.e., $\mathbf{H}_{\mathrm{BR}}^{\mathrm{LoS}}$, can be given as
@@
  • $\mathbf{a}_{\mathrm{R}}=\left[1,\cdots,e^{-j2\pi\frac{\Gamma^{\mathrm{br}}_{i_{z},i_{y}}}{\lambda}},\cdots,e^{-j2\pi\frac{\Gamma^{\mathrm{br}}_{I_{z},I_{y}}}{\lambda}}\right]^{T}$ denotes the steering vector of FD-RIS, where $\Gamma^{\mathrm{br}}_{i_{z},i_{y}}=(i_{z}-1)d\cos(\theta_{\mathrm{br}})+(i_{y}-1)d\sin(\theta_{\mathrm{br}})\cos(\varphi_{\mathrm{br}})$.
  • $\mathbf{a}_{\mathrm{B}}=\Big[1,\cdots,e^{-j2\pi\frac{(n_{\mathrm{t}}-1)d\sin(\theta_{\mathrm{br}})\sin(\varphi_{\mathrm{br}})}{\lambda}},\cdots,\\ e^{-j2\pi\frac{(N_{\mathrm{t}}-1)d\sin(\theta_{\mathrm{br}})\sin(\varphi_{\mathrm{br}})}{\lambda}}\Big]^{T}$ is the steering vector of the BS.
@@
$\mathbf{h}_{\mathrm{r}k}=\zeta(d_{\mathrm{r}k})\left(\sqrt{\frac{\beta_{2}}{1+\beta}}\mathbf{h}_{\mathrm{r}k}^{\mathrm{LoS}}+\sqrt{\frac{1}{1+\beta_{2}}}\mathbf{h}_{\mathrm{r}k}^{\mathrm{NLoS}}\right),$ (15)
where $\beta_{2}$ is the Rician factor of FD-RIS-users link, $\zeta(d_{\mathrm{r}k})$ denotes the large-scale path loss with $d_{\mathrm{r}k}$ being the distance between the FD-RIS and the $k$-th user. $\mathbf{h}_{\mathrm{r}k}^{\mathrm{NLoS}}$ is the NLoS component, and each element of $\mathbf{h}_{\mathrm{r}k}^{\mathrm{NLoS}}$ is i.i.d. and follows complex Gaussian distribution with zero mean and unit variance. $\mathbf{h}_{\mathrm{r}k}^{\mathrm{LoS}}$ represents the LoS component, which is given by
@@
$y_{k}=\mathbf{h}_{\mathrm{r}k}^{T}\tilde{\boldsymbol{\Theta}}\boldsymbol{\Theta}_{k}\mathbf{H}_{\mathrm{BR}}\mathbf{w}_{k}s_{k}+\sum_{j\neq k}^{K}\mathbf{h}_{\mathrm{r}k}^{T}\tilde{\boldsymbol{\Theta}}\boldsymbol{\Theta}_{k}\mathbf{H}_{\mathrm{BR}}\mathbf{w}_{j}s_{j}+n_{k},$ (17)
@@ III-B Problem Formulation
where $\omega_{k},~k\in\mathcal{K},$ is the weighted factor of the $k$-th user, which can be utilized to control the priority of the $k$-th user and the fairness among the users. $f_{\min}$ and $f_{\max}$ denote the lower and upper bounds of the modulation frequency ${f}_{l},$ for $l\in\mathcal{L}$. $P_{\mathrm{tmax}}$ is the maximum power budget of the BS. Actually, it is challenging to address this optimization problem, since the non-convexity of the objective function with respect to (w.r.t.) $\{\mathbf{w}_{k}\}_{k=1}^{K}$, $\mathbf{f}$ and $\boldsymbol{\tau}$. To handle this problem, the MMSE method [29] is first utilized to transform the objective function expressed in logarithmic form. Then, the alternating strategy is leveraged to divide the transformed optimization problem into three subproblems, which are respectively for the design of active beamforming, time-delay variables, and modulation frequency variables. Specifically, leveraging the MMSE method and introducing auxiliary variables $W_{k}$ and $\mu_{k}$, the communication rate $R_{k},~k\in\mathcal{K},$ can be converted as
@@
[…]
The analysis above indicates that the lower bound of $R_{k}$ in the $(q+1)$-th iteration can be obtained with the given $W_{k}^{(q)}$ and $u_{k}^{(q)}$. Note that the values of $W_{k}^{(q)}$ and $u_{k}^{(q)}$ are calculated by utilizing the optimal expressions in (21) with the achieved active beamforming, time delays and modulation frequency solutions in the $q$-th iteration. Hence, the lower bound of $R_{k}$ in the $(q+1)$-th iteration can be expressed as $R_{k}\geq\tilde{R}_{k}=\frac{\ln(W_{k}^{(q)})-W_{k}^{(q)}E_{k}^{(q)}+1}{\ln 2}.$ Thus, in $(q+1)$-th iteration, the original optimization problem can be transformed as
@@ IV-A Active beamforming design
 $\quad\text{ s.t. }~\sum_{k=1}^{K}\left\|\mathbf{w}_{k}\right\|^{2}\leq P_{\mathrm{tmax}}.$ (23)
@@ IV-B Passive Beamforming Design for Time Delay
From the expression of $\tilde{\theta}_{i}=e^{-j2\pi g{f}_{l}\tau_{i}},~i\in\mathscr{I}$, it is worth noting that regardless the value of ${f}_{l},l\in\mathcal{L}$, each element in the FD-RIS has the capability to induce a specific phase-shift on the reflected harmonic signal by introducing an appropriate time delay. Therefore, we can initially design the reflection coefficients generated by the time delays in a holistic manner. In other words, $\tilde{\theta}_{i}=e^{-j2\pi gf_{l}\tau_{i}},~i\in\mathscr{I},$ can be treated as the optimization variable and designed, rather than focusing separately on the design of $\tau_{i},~i\in\mathscr{I}$. Subsequently, it is straightforward to derive the corresponding time delays $\tau_{i},~i\in\mathscr{I}$ with given ${f}_{l},~l\in\mathcal{L}$ based on the obtained reflection coefficients $\tilde{\theta}_{i},~i\in\mathscr{I}$ in the final step of the iteration.
@@
 $~\text{ s.t. }~~|\tilde{\theta}_{i}^{*}|=1,~\angle\tilde{\theta}^{*}_{i}\in[0,2\pi],~\forall i\in\mathscr{I},$ (31)
@@
[…]
@@
where $\psi_{\mathrm{PR}}^{(e)}$ represents the Polak-Ribiere parameter [33] whose expression can be expressed as (35), shown at the top of this page. The operator $\langle\cdot,\cdot\rangle$ denotes the inner product operation.
@@
where $\mathscr{P}(\cdot)$ is the projection operator, which maps the updated $\boldsymbol{\vartheta}$ to the feasible region $\mathcal{S}$ by element-wise retraction, i.e., $\frac{(\boldsymbol{\vartheta})_{i}}{|(\boldsymbol{\vartheta})_{i}|},~i\in\mathscr{I}$. $\eta^{(e)}$ denotes the Armijo step in the $(e+1)$-th iteration. To obtain a suitable step size, the Armijo criteria [33] is usually leveraged, which is given by
$f\left(\mathscr{P}\left(\boldsymbol{\vartheta}^{(e)}+\eta^{(e)}\mathscr{D}^{(e)}\right)\right)$ $\geq f\left(\boldsymbol{\vartheta}^{(e)}\right)+$
$c_{1}\eta^{(e)}\operatorname{grad}f\left(\vartheta^{(e)}\right)^{H}\mathscr{D}^{(e)},$ (37)
where $c_{1}\in(0,1)$ is a parameter that controls the amount of decrease in the objective function. The details for solving the optimization problem (IV-B) utilizing the Riemannian conjugate gradient algorithm are presented in Algorithm 2.
@@
7: Output: the optimal $\boldsymbol{\vartheta}^{\mathrm{opt}}$.
@@ IV-C Passive Beamforming Design for Modulation Frequency
[…]
@@
  • $\rho^{(z)}=\frac{1}{10L}\operatorname{abs}\left(\frac{\partial\tilde{g}}{\partial\mathbf{f}}\left(\mathbf{f}^{(z-1)}\right)\right)^{T}\left(\mathbf{f}_{\max}-\mathbf{f}_{\min}\right)$ is the conservative factor, which can control the conservative property of the MMA convex approximation.
@@
In addition, the first-order gradients of the objective function w.r.t. the optimized variable, i.e., $\frac{\partial\tilde{g}}{\partial\mathbf{f}}$, can be derived as  [^4] As presented in Remark 4, during the alternating optimization process, the entire matrix $\tilde{\boldsymbol{\Theta}}$ can be treated as the optimization variable, without separately considering the modulation frequency variables contained within it. Thus, it can be treated as a known system parameter during the optimization of modulation frequencies.
@@
which represents the $l$-th element of $\frac{\partial\tilde{g}}{\partial\mathbf{f}}$. In addition, the $v$-th element of $\frac{\partial\boldsymbol{\vartheta}_{k}}{\partial{f}_{l}}$ is given as
  • $\mathcal{I}_{v}\left(\frac{\partial\boldsymbol{\vartheta}_{k}}{\partial{f}_{l}}\right)=\begin{cases}A_{0}j2\pi g\left(\frac{d^{\mathrm{r}k}_{v}}{c}-t\right)\times\\ e^{j2\pi g{f}_{l}\left(\frac{d^{\mathrm{r}k}_{v}}{c}-t\right)+\phi_{0}},&v\in\mathcal{V},\\ 0,&otherwise,\end{cases}$
where we have $\mathcal{V}\triangleq\{(r-1)SMN+\tilde{s}N+(m-1)SN+n\}_{m\in\mathcal{M},n\in\mathcal{N}}$ with $r=\lceil l/S\rceil,\tilde{s}=\operatorname{mod}(l-1,S)$. Note that $d^{\mathrm{r}k}_{v}$ denotes the distance between the $v$-th element of the FD-RIS and user $k$, which can be mapped as $d^{\mathrm{r}k}_{i_{z},i_{y}}$ easily.
@@
$\frac{\partial\widehat{g}}{\partial{f}_{l}}(\mathbf{f}|\mathbf{f}^{(z-1)})=\frac{p_{l}^{(z)}}{\left(u_{l}^{(z)}-{f}_{l}\right)^{2}}-\frac{q_{l}^{(z)}}{\left({f}_{l}-o_{l}^{(z)}\right)^{2}}.$ (43)
@@
In order to address this issue, an additional inner loop iteration is implemented to fine-tune the conservative factor $\rho$ for a more conservative convex approximation. This adjustment aims to satisfy the condition $\widehat{g}(\mathbf{f}^{(z)}|\mathbf{f}^{(z-1)})\geq\tilde{g}(\mathbf{f}^{(z)})$. Specifically, we can readily demonstrate that $\widehat{g}(\mathbf{f}|\mathbf{f}^{(z-1)})$ in (41) is a monotonically increasing function w.r.t. the conservative factor $\rho$. Therefore, if the solution of the MMA subproblem in the $z$-th outer-loop iteration does not satisfy the third condition, an inner-loop iteration is employed to progressively increase the conservative factor until the third condition is fulfilled. Fig. 3 illustrates how the conservative factor adjusts the conservativeness of the MMA convex approximation. In the $(x+1)$-th inner loop iteration, the update rule of the conservative factor can be given by
Figure 3: Schematic illustration of how the conservative factor $\rho$ adjusts the conservativeness of the MMA convex approximation.